OpenStudy (anonymous):

dy/dx=(1+2x^2)/(x cos y)

OpenStudy (anonymous):

What's the question?

OpenStudy (anonymous):

Find the general solution to the following differential equation :

myininaya (myininaya):

\[\cos(y) dy=\frac{1+2x^2}{x} dx\] integrate both sides

myininaya (myininaya):

\[\int\limits_{}^{}\cos(y) dy=\int\limits_{}^{}(\frac{1}{x}+2x) dx\]

myininaya (myininaya):

\[\sin(y)=\ln|x|+x^2+C\]

OpenStudy (anonymous):

∫ (1+2x^2)/(x cos y) dx=( X^2+log(x))sec(y)+c Is this correct myininaya ?

myininaya (myininaya):

we have to do separation of variables

myininaya (myininaya):

we had \[\frac{dy}{dx}=\frac{1+2x^2}{x \cos(y)}\] I want to get my y's together and my x's together so first thing i did was multiply cos(y) on both sides \[\cos(y) \frac{dy}{dx}=\frac{1+2x^2}{x}\] Then I multiplied dx on both sides \[\cos(y) dy=\frac{1+2x^2}{x} dx\]

myininaya (myininaya):

Now we can integrate both sides

OpenStudy (anonymous):

i have to say thank you as well.

myininaya (myininaya):

Aww... you're welcome :)